OpenStudy (anonymous):

solve the quadratic equation by completing the squares 2x^2 +4x+7=0 a. x=-2+/-sqrt /2 b. no real solution c. x=-2+/-sqrt14/2 d. x=+/- sqrt7

OpenStudy (anonymous):

No real solutions

OpenStudy (anonymous):

And you can do that by looking at the discriminant

OpenStudy (anonymous):

Do you know how to look at the discrimant and find whether it is real or complex?

OpenStudy (anonymous):

no i dont how is that done?

OpenStudy (anonymous):

Do you want to know, or you are satisfied with the answer

OpenStudy (anonymous):

lol id like to know please

OpenStudy (dumbcow):

factor out the 2 x^2 +2x +7/2 = 0 take half the second coefficient and square it (x+1)^2 -1 +7/2 = 0 (x+1)^2 = -5/2 x+1 = sqrt(-5/2) sqrt of neg num is non-real answer

OpenStudy (anonymous):

For any quadratic equation \[ax^2+bx+c\] if \[(b^2-4ac)\]<0 then the equation has no real roots

myininaya (myininaya):

lol thats blank above

OpenStudy (anonymous):

Just remember this and it will work always

OpenStudy (anonymous):

Did you understand that?

OpenStudy (anonymous):

any problem?

OpenStudy (anonymous):

no im reading it trying to understand thanks a bunch everyone :)

OpenStudy (anonymous):

yes i get it